prove that if 4x-4=3y+6 and y=2x then x=-5
by substitution y=2x so 4x-4 = 3(2x)+6
4x-4=6x+6
-4x from both sides -4=2x+6
-6 from both sides -10=2x
divide both sides by 2 -5=x

Found 2 solutions by drj, purplerain:
Answer by drj(1380)

prove that if 4x-4=3y+6 and y=2x then x=-5

Step 1. Substitute the values of x and y into the equation to see if it leads to a true statement.

Step 2. Find y when x=-5. Then y=2*(-5)=-10.

Step 3. Substitute x=-5 and y=-10 into 4x-4=3y+6

4*(-5)-4=3(-10)+6

-20-4=-30+6


-24=-24 which is a true statement.

Step 4. ANSWER: x=-5 and y=-10 satisfies both y=-2x and 4x-4=3y+6.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J



Answer by purplerain(1)

prove that if 4x-4=3y+6 and y=2x then x=-5
by substitution y=2x so, 4x-4 = 3(2x)+6
simplified expression becomes, 4x-4=6x+6
-4x from both sides, 4x-4x-4=6x-4x+6 equals -4=2x+6
-6 from both sides , -10=2x
divide both sides by 2, therefore -5=x or x=-5